OpenStudy (anonymous):

3000 is invested in an account that earns 2% per year in interest. if the interest is compounded monthly how long will i take until the account contains 5200.

OpenStudy (amistre64):

the rate per period is .02/12 the number of periods is, lets say, m for months take your initial deposit, and compound it for m periods to determine when it reaches the desired amount

OpenStudy (amistre64):

\[A=P(1+k)^m\] \[\frac{A}{P}=(1+k)^m\] \[log(\frac{A}{P})=m~log((1+k)\] \[\frac{log(\frac{A}{P})}{log(1+k)}=m\]